### v1

# Bridging Developer Instructions and Code Completion Through Instruction-Aware Fill-in-the-Middle Paradigm

Abstract.
Large Language Models (LLMs) have significantly advanced code completion, yet they often fail when the developer’s intent is underspecified in the code context. In practice, developers frequently add natural language instructions (e.g., comments) to clarify their intent. However, exisintg code LLMs applied for code completion systems merely undergoes a fill-in-the-middle (FIM) pre-training, which struggle to leverage this information effectively due to the lack of instruction-like training data. Existing instruction-tuning techniques, which improve instruction-following in general code generation, paradoxically degrade FIM performance, forcing a trade-off between instruction-following and infilling capabilities. To address this gap, we introduce Instruction-aware Fill-in-the-Middle (IFIM), an instruction-tuning method specifically designed to enhance FIM code completion models. IFIM extends the conventional FIM training objective by incorporating an explicit instruction section into the input, enabling the model to learn from (prefix, instruction, suffix) triplets. This approach allows the model to effectively leverage developer-provided directives while preserving its core completion abilities when no instructions are present. To facilitate this, we constructed a large-scale dataset by using GPT-4o to generate concise, intent-focused instructions for code infilling examples. We evaluated IFIM by applying it to two popular base models, Deepseek-Coder and Qwen2.5-Coder, on the benchmarks derived from HumanEval-infilling and RepoMasterEval. The results demonstrate that IFIM significantly improves instruction-following capabilities, boosting the Pass@1 score from 84.6% to 93.6% on HumanEval-infilling. Moreover, this enhancement does not compromise the models’ original performance on FIM code completion tasks with no instructions provided.
copyright: noneconference: 48th International Conference on Software Engineering; April 12–18, 2026; Rio De janeiro, Brazil
1. Introduction
Modern Integrated Development Environments (IDEs) rely on code completion to accelerate development (Wang et al., 2023): given a partial program, a model predicts the missing sections and displays them to the programmer. Large Language Models (LLMs) have pushed this feature from keyword-level suggestions to full-function generation, which adapts to different coding styles and languages with remarkable accuracy. As a result, coding assistants, with LLM-powered code completion as a key feature, have become a buzzword in the developer community, such as Cursor, Github Copilot, and Trea.
Figure 1. Guiding an LLM for code completion using instructive comments. Left: An ambiguous request leads to an unhelpful completion. Right: A specific comment clarifies the developer’s intent, resulting in the desired code.
Despite their capabilities, LLM-generated code completions are not always helpful to developers. Sun et al. (Sun et al., 2025) found that most of the unaccepted code completions are those produced from underspecified intent. As shown in Figure 1, the code context does not show clear intent, making it almost impossible for the LLMs, even for human experts, to infer the missing code. However, adding a simple clarifying comment like “#filter out negative numbers” transforms the task from almost impossible to trivial. We refer to natural language describing what the missing code is expected to achieve in the code context as code completion instructions. In practice, developers often, if not always, compose these instructions (e.g., as inline comments) to clarify their intent for LLM-based tools. A study on developers’ behavior reveals that programmers invest 11.56% of their coding time instructing these LLMs, a figure comparable to the 14.05% they spend writing the functionality itself (Mozannar et al., 2024). Thus, the ability to accurately follow developer instructions is paramount for a code completion system to provide helpful assistance.
An LLM can enhance its instruction-following ability through instruction-tuning (Wei et al., 2021): fine-tuning the LLM on (instruction, response) pairs so it learns to follow free-form natural language directives. It is now a standard training step for chat models, such as DeepSeek-R1 (Guo et al., 2025) and Qwen-3 (Yang et al., 2025). Instruction-tuned code LLMs, such as Magicoder (Wei et al., 2023) and Wizardcoder (Luo et al., 2023), can achieve substantial improvements in code generation (i.e., generating a complete function from a natural language description). However, it backfires for fill-in-the-middle (FIM) code completion (i.e., predicting the missing middle section given the prefix and suffix of the program). Code LLMs are found to exhibit a degradation in FIM code completion capabilities after the instruction-tuning training step (Pan et al., 2024; Ahmad et al., 2025; Wu et al., 2024). This is because current instruction-tuning methods for code LLMs are designed for generating complete, assistant-style outputs rather than context-aware code insertions. It leads to unpredictable behaviors, such as repeating the suffix or introducing unrelated code, in FIM code completion scenarios. To preserve FIM performance, industrial code completion systems like JetBrains’ Mellum (Pavlichenko et al., 2025) still rely on base pre-trained LLMs that are not instruction-tuned. Specifically, they are pre-trained with the FIM objective (Bavarian et al., 2022), which aims to grant them the capability to predict missing code within a given prefix and suffix. However, the public code repositories used for pre-training contain insufficient instruction-like data (Sun et al., 2022; Shi et al., 2022), which severely limits the base models’ ability to follow complex user directives.
Given the critical role of instruction following in code completion, discarding instruction tuning entirely seems unwise, especially considering its proven effectiveness in general code generation tasks. This presents an important research opportunity: Can we develop a specialized training method for instruction-aware code completion that addresses current limitations? An ideal solution should simultaneously: (1) enhance the model’s ability to comprehend and respond to developers’ code completion instructions, while (2) preserving (or even improving) the base model’s native code completion capabilities when no instructions are provided. To the best of our knowledge, this crucial gap remains unaddressed in both academic literature and industrial practice.
To fill this gap, we propose Instruction-aware Fill-in-the-Middle (IFIM), an instruction-tuning method specifically designed for FIM code completion. Rather than abandoning the FIM pretraining objective like existing instruction-tuning methods (Wei et al., 2023; Luo et al., 2023), IFIM extends it by injecting an instruction section into its input sequence. This simple but effective design preserves the full spectrum of the original FIM paradigm, ensuring that the model continues to perform gracefully when no instructions are present, while enabling LLMs to explicitly understand and follow developer-supplied instructions. Concretely, IFIM starts with constructing an instruction dataset by decomposing source code into prefix, middle, suffix triplets and prompting an LLM like GPT-4o to emit a concise one-sentence description of the middle segment’s intent as the instruction. The resulting instruction dataset is then used to train the LLMs with the IFIM objective, which enforces the model to predict the middle segment given the prefix and suffix, followed by the instruction. It thereby improves the model’s ability to generate contextually coherent and instruction-aligned infilling code.
We evaluate IFIM on two code completion benchmarks IHumaneval and IRepoMasterEval, derived respectively from Humaneval-infilling (Fried et al., 2022) and RepoMasterEval (Wu et al., 2024) with code completion instructions added to their data samples. To investigate the impact of IFIM training on widely used code completion LLMs, we experiment with two base code LLMs: Deepseek-Coder (DeepSeek-AI et al., 2024) and Qwen2.5-Coder (Hui et al., 2024). Our comprehensive experiments demonstrate that IFIM achieves remarkable improvements in instruction-following capability while maintaining the LLMs’ ability to complete the code without instructions. Specifically, given the same instructions, our IFIM-trained variants can achieve better performance than the base model, with Pass@1 score averagely improved from 87.8% to 94.7% on IHumaneval and from 14.7% to 20.7% on IRepoMasterEval. Crucially, when no instructions are provided, IFIM models either match or slightly exceed their base counterparts’ performance. For example, on the IHumaneval benchmark, IFIM improves Deepseek-Coder from 68.6% to 78.2% and Qwen2.5-Coder from 76.0% to 76.3%. Moreover, we also conducted a series of ablation experiments to analyze the optimal position for inserting the instruction component, the ratio of IFIM processed data, and the advantages of using a standalone instruction component. These results position IFIM as a practical, backward-compatible upgrade path for existing LLM-based code completion systems.
Our contributions can be summarized as follows:
  • We identify and address a critical conflict between instruction-following and fill-in-the-middle capabilities in existing code completion LLMs. We propose a simple but effective solution, Instruction-aware Fill-in-the-Middle (IFIM), an instruction-tuning method that enables FIM-pretrained models to better follow natural language instructions without degrading their performance for non-instructed requests.
  • We construct and release a large-scale dataset for instruction-aware code completion. This dataset, containing 122.9k multi-language data samples, provides a valuable resource for training and evaluating models on their ability to follow user intent during infilling tasks.
  • We conduct extensive experiments demonstrating IFIM’s dual benefit. Compared to the base models, IFIM significantly enhances instruction-following, boosting the average Pass@1 score from 87.8% to 94.7% on HumanEval-infilling and from 14.7% to 20.7% on RepoMasterEval. Crucially, it achieves this while preserving or even improving performance when no instructions are given, validating IFIM as a practical and effective method for building more helpful code completion systems.
Our source code and datasets are released at https://anonymous.4open.science/r/ifim-artifacts.
2. Background and Motivation
In this section, we briefly recap the Fill-in-the-Middle training objective and pinpoint the complementary weaknesses that keep pre-trained code LLMs and instruction-tuned code LLMs from fully covering real-world developer needs.
2.1. Fill-in-the-Middle Code Completion
The Fill-in-the-Middle (FIM) objective (Bavarian et al., 2022) was proposed to extend the conventional left-to-right, causal language-modeling paradigm with a crucial infilling capacity during pre-training. It grants the model the ability to generate a contiguous span of text that is missing from the middle of a document while conditioning on both the preceding (prefix) and the succeeding (suffix) context. Code completion is the natural test-bed for this capability, as developers routinely need to insert new logic between already-written lines, such as the body of a newly declared function, missing import statements, or boilerplate loops. An illustrative example of this is shown in Figure 1. Because the content that follows the developer’s cursor is already known, a model that can exploit this suffix context is strictly more informative and powerful than one restricted to a left-only prefix.
Standard causal language modeling represents the document as a sequence of tokens, following the left-to-right order. In contrast, FIM splits a training sequence into prefix $P$, middle $M$, and suffix $S$, then rearranges them with special tokens, e.g.,
$\texttt{<PRE>}\,\circ\,P\,\circ\,\texttt{<SUF>}\,\circ\,S\,\circ\,\texttt{<MID>}$
where $\circ$ denotes concatenation and <PRE>, <SUF>, and <MID> are the special tokens respectively for the prefix, suffix, and middle sections. For example, a piece of code like ‘for i in range(10): print(i)’ could be partitioned into $P$: ‘for i in ’, $S$: ‘ print(i)’, and $M$: ‘range(10):’. The model would then see the transformed input ‘¡PRE¿for i in ¡SUF¿ print(i)¡MID¿’. The training objective is to maximize the log-likelihood of the middle span $M$, i.e, the ‘range(10)’ in this example. This data transformation maintains the model’s underlying architecture but forces it to learn code completion with bidirectional information. Crucially, this approach does not sacrifice the model’s standard left-to-right language modeling ability. By training on a carefully balanced mix of regular and FIM-transformed samples, the model retains its next-token prediction performance while gaining the powerful new infilling skill. In practice, this technique has been widely adopted, and nearly all recent state-of-the-art code LLMs, such as DeepSeek-Coder-v2 (DeepSeek-AI et al., 2024), CodeGemma (Team et al., 2024), Seed-Coder (Zhang et al., 2025), and Qwen2.5-Coder (Hui et al., 2024), are pretrained with the FIM objective.
2.2. Base Code LLMs are Sub-Optimal Instruction Followers
Base code LLMs are pre-trained on a vast data corpus, the vast majority of which consists of code files scraped from public repositories, such as GitHub. Consequently, during code completion, they fill the gap with whatever appears most likely based on patterns learned from the training data. This extends to natural language instructions written as comments; the model “follows” them by predicting the most probable code sequence found after similar comments in its training data. However, the developer’s true, fine-grained intent while coding is a signal that is largely absent from the public code repositories used for pre-training. Much of that intent never reaches the committed source: it may surface as a fleeting search-engine query, a short Slack thread with a teammate, or even remain unspoken in the developer’s mind. The comments that are finally documented in the code are, therefore, an insufficient proxy for instructions. Moreover, prior studies (Sun et al., 2022; Shi et al., 2022) have shown that the comments in source code are very noisy, with only a small fraction of comments actually describing the functionality of the code. The majority serve other purposes, such as licensing information, TODOs, or internal team coordination, making them poor indicators of intent. Relying on them alone leaves the model under-conditioned for real-world scenarios where intent is expressed ad-hoc. To close the gap, we need to supplement the model’s strong FIM foundation with targeted instruction alignment that empowers it to interpret explicit natural-language instructions.
2.3. Conversational LLMs are Sub-Optimal Code Completers
Conversational LLMs, such as GPT-4o, Claude-3, and DeepSeek-R1, have achieved remarkable performance on code-related dialog tasks. These include answering debugging questions, generating entire programs from natural-language specifications, and refactoring code snippets. In these scenarios, the model receives a textual instruction and produces a helpful, conversational response that mixes text and code. This capability is cultivated through an instruction tuning stage, where models learn to follow human instructions and dialog formats, often through supervised learning on instruction-response pairs (Wei et al., 2021). However, the datasets used for this purpose, such as Magicoder (Wei et al., 2023) and Wizardcoder (Luo et al., 2023), primarily emphasize generating complete, assistant-style outputs rather than context-aware code insertions. This focus comes at a cost: the training process largely overlooks the strict contextual and logical consistency required by FIM code completion, where new code must integrate seamlessly with its surrounding prefix and suffix. As a result, when applied to FIM tasks, these powerful conversational models exhibit various failure modes, such as repeating the suffix or generating irrelevant content (Ahmad et al., 2025; Gong et al., 2024; Pan et al., 2024). An FIM-trained base model like Codegemma-7b (Team et al., 2024) can easily outperform the much larger conversational LLM, GPT-4o, on real-world code completion benchmarks (Pan et al., 2024). Consequently, the LLMs deployed in production code completion systems are typically pre-trained base models, rather than their instruction-tuned counterparts.
3. Instruction-aware Fill-In-the-Middle
Figure 2. The overall framework of IFIM, which consists of synthesizing an instruction dataset and performing an instruction tuning phase.
In this section, we describe our Instruction-aware Fill-In-the-Middle (IFIM) method. As illustrated in Figure 2, IFIM augments a FIM-trained code LLM with instruction-following capabilities through a two-step process: (1) synthesizing a specialized dataset for instruction-guided code completion, and (2) performing an instruction tuning phase designed to learn from the instructions while preserving the model’s original infilling performance.
3.1. Instruction Synthesis for Code Completion
Our approach begins by creating an instruction dataset tailored for code completion. The synthesis process starts with code sourced from open-source repositories. For each code sample, we randomly select a contiguous block of one or more lines to serve as the middle segment, with the surrounding code forming the prefix and suffix. Following established practices for creating code instruction datasets (Luo et al., 2023; Wei et al., 2023; Huang et al., 2024), we then prompt a powerful general-purpose LLM (e.g., GPT-4o) to generate a corresponding code completion instruction based on all three segments.
Specifically, we present the entire code sample to the LLM, but wrap the middle part with ¡explain¿¡/explain¿ tags and request a concise, one-sentence functional description. The prompt is structured as follows:
[…]
The LLM’s response serves as the instruction, which we combine with the code segments to form a (prefix, middle, suffix, instruction) tuple. Finally, we perform two data filtering steps. To ensure instructions are concise, we decontaminate the dataset by discarding any generated instructions longer than a single sentence. While some of the remaining instructions may have minor imperfections (e.g., being overly brief or vague), we retain them as they reflect realistic scenarios developers often face. While model-generated instructions may contain some imperfections (such as brevity or vagueness), we retain them as they represent realistic instruction scenarios developers encounter. Additionally, following (Huang et al., 2024), we exclude any samples that match problems from common evaluation benchmarks, such as HumanEval (Chen et al., 2021) and MBPP (Austin et al., 2021).
3.2. Instruction Tuning after FIM-based Pretraining
Our instruction tuning process builds upon a FIM-pretrained LLM, extending its capabilities to understand natural language instructions while carefully preserving its foundational infilling skills. Each training sample consists of four components: prefix ($P$), middle ($M$), suffix ($S$), and instruction ($I$). The model learns to predict the middle segment $M$ conditioned on the combined context of $P$, $S$, and $I$. This forms a supervised fine-tuning task where the objective is to maximize the log-likelihood of the target sequence:
(1) $\max_{\theta}\mathbb{E}{(P,M,S,I)\sim\mathcal{D}}\left[\log P_{\theta}(M|P,S,I)\right]$
where $\mathcal{D}$ is our synthesized dataset, and the input is formatted according to a specific ordering of the components.
3.2.1. IFIM Mode
A critical design choice for IFIM is the sequential ordering of the four components ($P$,$S$,$M$,II) in the input sequence, which we term the IFIM mode. To maintain the model’s pretrained abilities, we preserve the relative ordering of the original FIM components ($P$,$S$,$M$) and insert the instruction $I$. This approach avoids the performance degradation often seen when existing instruction tuning methods are applied to FIM models.
For any given FIM pretraining mode, such as PSM (Prefix-Suffix-Middle), there are four potential IFIM modes (e.g., IPSM, PISM, PSIM, PSMI). Through comprehensive ablation studies (detailed in Section 5.2), we empirically found that placing the instruction immediately before the middle component consistently yields the best performance. Therefore, we adopt this strategy as our default. For the three most common FIM modes, PSM, PMS, and SPM, our corresponding default IFIM modes are PSIM, PIMS, and SPIM.
Notably, the instruction’s position can impact the efficiency of the key-value cache during inference. Placing the instruction at the beginning of the input (e.g., IPSM) means that any minor change to the instruction would invalidate the cache for all subsequent tokens, leading to wasted computation. This is not a concern for PSIM and SPIM modes, as the instruction appears late in the sequence. For models using the PMS mode, the corresponding PIMS mode would invalidate the suffix’s cache. In such cases, a PMSI configuration may offer a better trade-off between performance and caching efficiency.
3.2.2. Training Sample Construction
With the IFIM modes defined, we formalize the construction of input-output pairs for instruction tuning. The ground-truth output is simply the middle segment $M$. The input, however, must be carefully formatted. Let’s use the PIMS mode as an example. Its underlying FIM format (PMS) is:
$\texttt{<PRE>}\,\circ\,P\,\circ\,\texttt{<MID>}\,\circ\,S\,\circ\,\texttt{<SUF>}\hfill$
where $\circ$ denotes concatenation and <PRE>, <SUF>, and <MID> are the special tokens respectively for the prefix, suffix, and middle sections. For IFIM, we insert the instruction $I$ at the designated position, i.e, between the prefix $P$ and the <MID> token for PIMS. To delimit the inserted $I$ from its adjacent code segments, i.e., the $P$ in this case, we insert a new special token <INS> to serve as the delimiter. Therefore, the input of the training sample in PIMS mode will be constructed as follows:
$\texttt{<PRE>}\,\circ\,P\,\circ\,\texttt{<INS>}\,\circ\,I\circ\,\texttt{<MID>}\circ\,S\,\circ\,\texttt{<SUF>}\hfill$
This construction strategy applies similarly to other modes. For example, the SPIM and PSIM modes are constructed as follows:
$\texttt{<SUF>}\,\circ\,S\,\circ\,\texttt{<PRE>}\,\circ\,P\,\circ\,\texttt{<INS>}\,\circ\,I\,\circ\,\texttt{<MID>}\hfill$
$\texttt{<PRE>}\,\circ\,P\,\circ\,\texttt{<SUF>}\,\circ\,S\,\circ\,\texttt{<INS>}\,\circ\,I\circ\,\texttt{<MID>}\hfill$
Adding a new token to an LLM’s vocabulary would typically require reinitializing the language model head, risking knowledge forgetting. To avoid this, we repurpose an existing but rare token from the vocabulary (e.g., the one with the lowest frequency) to serve as our <INS> token.
Figure 3. A usage example of IFIM-trained code LLMs, where the code context is first propcessed into IFIM format and then infilled by an IFIM-trained code LLM.
3.3. Usage for Code Completion
An IFIM-trained LLM enables a developer to guide code completion by embedding instructions directly within their code. Since plain text would violate the programming language’s syntax, we propose writing instructions as specially marked inline comments. An example of this is shown in Figure 3. In this example, a developer using Python could write “#!filter out negative numbers” as the instruction. The ‘#!’ prefix acts as a unique marker that distinguishes an IFIM instruction from a regular comment. This allows client-side tools, like IDE plugins, to build user-friendly features around it. For example, an IDE could be configured to automatically remove the instructional comment once the generated code is accepted by the user. When a code completion request is made, the code completion system parses the code around the cursor to extract the prefix and suffix, and identifies the specially marked comment as the instruction. These three components, prefix, suffix, and instruction, are then formatted into the required IFIM input sequence. The IFIM-trained LLM receives this input and generates the code to fill the middle, which is then presented to the developer as a completion suggestion.
4. Experiment Setup
To demonstrate the effectiveness of our proposed method, we conduct a comprehensive evaluation across multiple dimensions. This section presents the experimental configuration, including research questions, instruction dataset construction, baseline models, evaluation benchmarks, and implementation details. Our experimental design is structured to address the following four research questions:
  • RQ1: How effective is IFIM in improving the instruction following capability while maintaining the infilling capability of code LLMs for code completion?
  • RQ2: How do different IFIM modes affect the performance of IFIM-trained code LLMs?
  • RQ3: What is the impact of varying the proportion of IFIM-processed data in the instruction tuning dataset?
  • RQ4: Does representing instructions as a dedicated component outperform using instructions as inline comments?
4.1. Instruction Dataset
Following the instruction synthesis process in Section 3.1, we construct a comprehensive instruction dataset for our experiments. The code snippets are sourced from OSS-Instruct (Wei et al., 2023) and evol-codealpaca (Luo et al., 2023), two well-established open-source instruction datasets. For each data point, we retain the code samples while removing the original instruction components, yielding 122.9k multi-language samples with 70.0% written in Python. For each code snippet, we randomly select 1-3 contiguous lines as the middle span and employ OpenAI GPT-4.1 (gpt-4.1-2025-04-14) as the instruction synthesizer to generate corresponding natural language descriptions.
4.2. Code LLMs
We evaluate our approach using two popular code LLMs as base models: Deepseek-Coder (Guo et al., 2024) and Qwen-2.5-Coder (Hui et al., 2024). Each model family includes a base version (pretrained with FIM objectives) and an instruction-tuned version (further fine-tuned on proprietary instruction datasets). In our experiments, we apply IFIM training to the base versions since they are the LLMs used for code completion.
  • Deepseek-Coder (6.7B): Deepseek-Coder is a specialized code language model designed to excel in FIM tasks. It is commonly adopted as the baseline for various studies.
  • Qwen-2.5-Coder (7B): Qwen-2.5-Coder is an iteration of Alibaba’s Qwen series, optimized for code-related tasks.
4.3. Evaluation Benchmarks
We assess the instruction-following capabilities of our models using benchmarks derived from two complementary ones: IHumanEval from HumanEval-infilling and IRepoMasterEval (IRME) from RepoMasterEval, which respectively represent controlled and real-world evaluation scenarios.
  • HumanEval-infilling (Fried et al., 2022) is a widely-used FIM benchmark constructed by removing middle spans from the canonical solutions in HumanEval. We adopt its single-line version, consisting of 1,640 infilling tasks. Each task includes a Python function with one random line of missing code, test cases, and docstrings describing the function’s functionality. While widely adopted, this benchmark has notable limitations for real-world applicability: its function-level scope and well-documented examples differ from practical code completion scenarios. However, considering its popularity, we still adopt it as one evaluation benchmark for our experiments. To better simulate real-world conditions and increase the code completion difficulty, we remove the docstrings from all evaluation samples.
  • RepoMasterEval (RME) (Wu et al., 2024) provides a more realistic evaluation framework for code completion models, which is used by ByteDance to evaluate their commercial code completion models. The pass rate on this benchmark is found to be highly correlated with the online user acceptance rate. Unlike HumanEval-infilling’s function-level focus, RepoMasterEval is built upon real-world code repositories. It extracts the code files from various repositories and processes them into code infilling tasks, i.e., code files with missing middle spans. The model outputs are evaluated using the unit tests of the repository, enhanced through mutation testing and manual case additions.
    Due to computational constraints limiting our models’ context windows, we retain the last 20 lines of the prefix and the first 20 lines of the suffix for each sample.
Notably, these two benchmarks are not natively paired with code completion instructions. To evaluate the instruction-following capability of the models, we synthesize instructions for each sample using OpenAI GPT-4.1 (gpt-4.1-2025-04-14) with the same prompt employed in our instruction synthesis process. To further ensure the quality of the instructions, we manually revise those of low quality. Given the large sample sizes, we randomly select statistically representative subsets from each benchmark (95% confidence level, 5% margin of error, 50% population proportion), resulting in 312 and 256 samples for HumanEval-infilling and RepoMasterEval, respectively. The manual filtering process involves two authors independently reviewing all generated instructions and revising those that inadequately describe the desired code functionality. The authors achieve an inter-rater agreement of 0.72 and collaboratively resolve disagreements through discussion. The final curated instruction benchmarks are respectively named as IHumanEval and IRepoMasterEval (IRME), which are used in the experiments.
4.4. Implementation Details
In our experiments, we use the Huggingface Transformers (Wolf et al., 2020) library with PyTorch to implement the models. The models are trained on a machine with 128 vCPUs, 200GB RAM, and two RTX A6000 GPUs (48GB RAM). For Deepseek-Coder, we adopt a batch size 2 with 128 accumulation steps and 1216 context length. For Qwen2.5-Coder, we adopt a batch size 1 with 256 accumulation steps to fit our GPU memory. These models are trained with an Adafactor optimizer (Shazeer and Stern, 2018) for 2 epochs. We set the initial learning rate at 5e-5 with 15 warmup steps and a linear scheduler. During inference, we use the greedy decoding strategy, and a maximum of 128 tokens are generated. For FIM mode selection, we adopt the recommended FIM mode based on the default usage example in their official documents: PMS for Deepseek-Coder and PSM for Qwen2.5-Coder.
5. Results
In this section, we report our experimental results and answer the four research questions.
5.1. RQ1: Effectiveness of IFIM
In this RQ, we systematically evaluate the effectiveness of IFIM, where we aim to assess two aspects: (1) whether IFIM can improve the instruction-following ability of the LLMs, and (2) whether IFIM can maintain the code completion ability when no instruction is given. Specifically, for the two base LLMs, Deepseek-coder and Qwen2.5-coder, we conduct IFIM training and obtain corresponding IFIM variants. The obtained models are evaluated with the two benchmarks, IHumanEval and IRME, under two settings: (1) with instruction and (2) without instruction. In the foIRMEr setting, we pair each completion request with its corresponding instruction, where the instructions for base models are written as a comment appended to the end of the prefix, while the instructions for IFIM variants are constructed following the IFIM mode. In the latter setting, we use pure prefixes and suffixes without any modification. The Pass@1 score is computed to assess the quality of the completion.
Table 1. Evaluating IFIM Effectiveness: A Performance Comparison with Base Models. The setting w/ ins. for base model means that we append the instruction as a comment to the prefix.
Setting Model Version Mode IHumanEval IRME
w/ ins. Deepseek-Coder base PMS 84.6% 10.9%
IFIM PIMS 93.6% 21.1%
Qwen2.5-Coder base PSM 91.0% 18.4%
IFIM PSIM 95.8% 20.3%
w/o ins. Deepseek-Coder base PMS 68.6% 7.4%
IFIM PIMS 78.2% 16.0%
Qwen2.5-Coder base PSM 76.0% 10.2%
IFIM PSIM 76.3% 13.3%
As reported in Table 1, the IFIM variants consistently outperform the base models when instructions are provided, across all the models and benchmarks. Specifically, on the IHumanEval benchmark, the Pass@1 scores after IFIM training improved from 84.6% to 93.6% for Deepseek-Coder and from 91.0% to 95.8% for Qwen2.5-Coder. A similar and even more pronounced trend is observed on the IRME benchmark, where Deepseek-Coder’s performance nearly doubled from 10.9% to 21.1%, and Qwen2.5-Coder’s score increased from 18.4% to 20.3%. This clearly suggests an enhanced instruction-following ability after IFIM training.
In the meantime, when no instructions are provided, the IFIM variants still maintain a comparable, or even better, performance than the base models. On IHumanEval, the performance is largely maintained, with Deepseek-Coder improving from 68.6% to 78.2% and Qwen2.5-Coder showing a slight gain from 76.0% to 76.3%. On the IRME benchmark, the IFIM variants again show notable improvements, with Deepseek-Coder increasing from 7.4% to 16.0% and Qwen2.5-Coder rising from 10.2% to 13.3%. It demonstrates that IFIM not only preserves but can also boost the infilling capability acquired during FIM pre-training. Given the above results, we conclude that IFIM is an effective and feasible method for code LLMs to enhance their overall performance.
Answer to RQ1: Training with IFIM substantially improves instruction-following ability. On average, it boosts performance with a relative increase of 8.0% on IHumanEval and a remarkable 52.0% on IRME when instructions are provided. Crucially, IFIM also enhances the original infilling capabilities. When no instructions are given, models show an average relative performance growth of 40.3% across both benchmarks.
5.2. RQ2: Impact of IFIM Mode
For RQ2, we aim to investigate the impact of different IFIM modes on the instruction-following ability of IFIM-trained LLMs. Specifically, for each base Code LLM, we conduct IFIM training with all its possible IFIM modes: IPMS, PIMS, PMIS, and PMSI for Deepseek-Coder; IPSM, PISM, PSIM, and PSMI for Qwen2.5-Coder, following the IFIM mode construction method in Section 3.1. In total, we have two base models and eight IFIM variants. Similar to RQ1, we evaluate the performance of these IFIM variants on IHumanEval and IRME, with and without instructions in the inputs, and compute the Pass@1 scores to measure the correctness of the completion.
Table 2. Impact of IFIM Mode on Instruction-Following Performance
Setting Model Version Mode IHumanEval IRME
w/ ins. Deepseek-Coder base PMS 84.6% 10.9%
IFIM IPMS 92.6% 19.9%
PIMS 93.6% 21.1%
PMIS 91.6% 20.3%
PMSI 92.3% 19.1%
Qwen2.5-Coder base PSM 91.0% 18.4%
IFIM IPSM 78.2% 12.9%
PISM 95.2% 19.5%
PSIM 95.8% 20.3%
PSMI 93.6% 18.0%
w/o ins. Deepseek-Coder base PMS 68.6% 7.4%
IFIM IPMS 77.9% 12.9%
PIMS 76.6% 16.0%
PMIS 78.2% 12.1%
PMSI 77.6% 13.7%
Qwen2.5-Coder base PSM 76.0% 10.2%
IFIM IPSM 60.3% 9.0%
PISM 76.3% 13.7%
PSIM 76.3% 13.3%
PSMI 76.6% 14.1%
We report the results in Table 2. The results reveal that the instruction-following ability acquired from IFIM training is sensitive to the position of the Instruction component. For example, the IFIM variant of Qwen2.5-Coder in PSIM mode achieves a 95.8% Pass@1 score on IHumanEval and 20.3% on IRME, but these scores plummet to 78.2% and 12.9% respectively in IPSM mode, falling below the base model’s performance.
Although performance varies across different modes, IFIM training leads to an overall improvement in instruction-following ability while generally preserving the model’s original infilling capability. Specifically, with the notable exception of the IPSM mode for Qwen2.5-Coder, all other seven IFIM variants outperform their respective base models on both benchmarks when instructions are provided. Even without instructions, most variants maintain or enhance performance over the base model, confirming the broad effectiveness of IFIM.
Notably, the I-before-M modes (i.e., placing the Instruction component immediately before the Middle component) consistently achieve the highest instruction-following performance compared with other IFIM modes. For Deepseek-Coder, the PIMS mode is the top perfoIRMEr on both IHumanEval (93.6%) and IRME (21.1%). Similarly, for Qwen2.5-Coder, the PSIM mode achieves the best scores on both IHumanEval (95%) and IRME (20.3%). This demonstrates that placing the instruction immediately before the code to be generated most effectively enhances the LLMs’ understanding and execution of the given instructions. We therefore adopt I-before-M as the default mode for IFIM training.
Answer to RQ2: IFIM mode has a non-trivial impact on the instruction-following ability of IFIM-trained LLMs. The I-before-M modes (placing Instruction before Middle) are demonstrably superior, outperforming the average of other modes by 4.1 percentage points on IHumanEval and 2.4 percentage points on IRME.
Table 3. Impact of Instruction Data Ratio in Training on Model Performance
Setting Model Mode Ratio IHumanEval IRME
w/ ins. Deepseek-Coder PMS base 84.6% 10.9%
PIMS 0% 84.6% 10.5%
25% 86.2% 18.0%
50% 90.1% 19.1%
75% 91.3% 19.1%
100% 93.6% 21.1%
Qwen2.5-Coder PSM base 91.0% 18.4%
PSIM 0% 90.8% 0.4%
25% 94.9% 20.3%
50% 73.4% 18.4%
75% 94.9% 18.8%
100% 95.2% 20.3%
w/o ins. Deepseek-Coder PSM base 70.7% 7.4%
PIMS 0% 72.4% 7.0%
25% 77.9% 12.9%
50% 77.6% 14.1%
75% 78.5% 12.1%
100% 78.2% 16.0%
Qwen2.5-Coder PSM base 76.0% 10.2%
PSIM 0% 76.2% 0.8%
25% 79.2% 14.8%
50% 55.8% 12.9%
75% 77.2% 12.9%
100% 76.3% 13.3%
5.3. RQ3: Impact of the Ratio of Instruction Data
In this section, we investigate how the proportion of instruction-formatted data in the training set affects model performance. Our goal is to find the ideal balance for our IFIM training. We experimented with five different data mixtures for our training set, converting 0%, 25%, 50%, 75%, and 100% of the data into our IFIM format. The remaining data in each mixture was kept in the standard FIM format native to the base models. For this experiment, we used the best-performing IFIM mode for each model: PIMS for Deepseek-Coder and PSIM for Qwen2.5-Coder. The 0% ratio serves as a crucial control group. In this case, the models were trained only on the code snippets from our dataset without our added instructions. This allows us to isolate the impact of the instructions themselves versus the impact of simply training on more code. Each of these five datasets was used to fine-tune a base model, and we evaluated their code completion accuracy (Pass@1) on the IHumanEval and IRME benchmarks, both with and without instructions.
The results in Table 3 reveal several interesting findings about the impact of instruction data ratio. First, we observe that increasing the ratio generally leads to better instruction-following performance. For both Deepseek-Coder and Qwen2.5-Coder, the peak Pass@1 scores are achieved with a 100% instruction data ratio. This holds true not only on IHumanEval (93.6% and 95.2%) but also on the IRME benchmark (21.1% and 20.3%). This suggests that to maximize a model’s ability to understand and follow instructions, a full conversion of the training data to the IFIM format is the most effective strategy. Second, the relationship is not strictly linear and reveals complex trade-offs. The most striking example remains the Qwen2.5-Coder, where the 50% ratio causes a significant performance drop on IHumanEval (73.4%) compared to both lower and higher ratios. This nonlinearity implies that the training process is sensitive to the specific balance between learning new instruction-based patterns and reinforcing existing ones.
For tasks without instructions where the model must rely on its original FIM capabilities, the optimal strategy is more nuanced. For Qwen2.5-Coder, a 25% ratio is consistently superior, achieving the best scores on both IHumanEval (79.2%) and IRME (14.8%). For Deepseek-Coder, the ideal ratio depends on the benchmark: a 75% ratio is best for IHumanEval (78.5%), while a 100% ratio is surprisingly optimal for IRME (16.0%). This demonstrates that the ideal mixture can be model-dependent and task-dependent.
Notably, the 0% ratio control group provides a crucial insight. When evaluated with instructions, this group shows a catastrophic degradation of performance on the IRME benchmark, with Qwen2.5-Coder’s score collapsing to 0.4%. Even Deepseek-Coder’s performance slightly degrades. This result highlights that simply training on more in-domain code (function-level snippets) without instructions is not only ineffective for instruction-following but can be actively harmful to the model’s ability to generalize to the real-world code distributions in IRME. The introduction of instructions, even with the same underlying code, mitigates this degradation. This suggests that the IFIM training method does more than just enhance instruction-following; it also appears to improve the model’s generalization, making it more robust against distribution shifts. A further investigation into this unexpected benefit of IFIM could be an important area for future work.
Answer to RQ3: For maximizing instruction-following ability, a 100% IFIM data ratio is optimal, boosting Pass@1 scores by up to 9.0 percentage points on IHumanEval and 10.2 points on IRME. Critically, the performance gains are unequivocally from the instructions themselves; training on just more code (the 0% ratio) leads to catastrophic degradation on out-of-distribution tasks, causing the IRME Pass@1 score to plummet for Qwen2.5-Coder (from 18.4% to 0.4%) when tested with instructions.
5.4. RQ4: Impact of Standalone Instruction Component
In this RQ, we seek to investigate the impact of our newly introduced Instruction component, compared with simply using instruction as inline comments. Specifically, we derive two training datasets from our instruction dataset: one processed using the IFIM format and the other by appending the instruction as an inline comment to the prefix. Using these two derived datasets, we train each base model (Deepseek-Coder and Qwen2.5-Coder) and correspondingly obtain two variants. Finally, we have two base models, each with one IFIM variant and one experimental variant for inline comments (named as CFIM, short for Comment-trained FIM). Since the instruction is appended as a part of the prefix, the inference for CFIM variants follows the same mode as their base models. We evaluate the performance of these models on IHumanEval-infilling and RepoMasterEval. During evaluation, the IFIM models are provided with instructed inputs, and the base model and CFIM models are provided with commented inputs (i.e., instructions are appended to the prefix as an inline comment). We compute the Pass@1 scores to measure the correctness of the completion.
Table 4. IFIM vs. Commented FIM (CFIM): Evaluating the Standalone Instruction Component
Model Version Mode IHumanEval IRME
Deepseek-Coder base PMS 84.6% 10.9%
CFIM PMS 79.0% 4.3%
IFIM PIMS 95.9% 21.1%
Qwen2.5-Coder base PSM 89.4% 18.4%
CFIM PSM 88.5% 18.8%
IFIM PSIM 96.6% 20.3%
The results presented in Table 4 reveal that simply adding instructions as inline comments is an ineffective and often detrimental strategy. For Deepseek-Coder, this approach results in degraded performance, with Pass@1 scores dropping from 84.6% to 79.0% on IHumanEval and from 10.9% to a mere 4.3% on IRME. The Qwen2.5-Coder CFIM variant showed a similar, though less pronounced, negative effect, with performance decreasing on IHumanEval and remaining stagnant on IRME.
This confirms a hypothesis: LLMs can be confused when instructional text is mixed with code as comments, negatively impacting their primary code completion capabilities. The stark contrast between the significant gains from IFIM and the performance degradation from CFIM suggests that the structural separation of instructions from the code, as implemented in IFIM, is essential for the model to properly disambiguate, understand, and benefit from the provided guidance.
Answer to RQ4: The structural separation of instructions from prefix and suffix better facilitates LLMs’ understanding of natural language instructions during training.
6. Related Work
6.1. Code LLMs for Code Completion
Code LLMs have become foundational in automated code completion. Early models like Codex (Chen et al., 2021), CodeGen (Nijkamp et al., 2022), and CodeGeex (Zheng et al., 2023) relied on left-to-right (L2R) generation, limiting their ability to handle real-world editing scenarios where code infilling is required. This gap led to the emergence of FIM training objectives (Bavarian et al., 2022; Fried et al., 2022), now adopted by state-of-the-art models including CodeGemma (Team et al., 2024), DeepSeek-Coder-V2 (DeepSeek-AI et al., 2024), and Qwen3-Coder (Yang et al., 2025). Various studies (Sagtani et al., 2025; Gong et al., 2025; Ding et al., 2024) are proposed to enhance FIM. For example, Sagtani et al. (Sagtani et al., 2025) introduce curriculum learning and context-aware sampling strategies to enhance FIM-based completion. Gong et al. (Gong et al., 2025) incorporate structural information (e.g., AST-guided masking) into the FIM pretraining process to better align the model with syntactic code editing. Ding et al. (Ding et al., 2024) propose a horizon-length prediction objective to dynamically guide the model’s infilling scope. Despite these enhancements, both standard and improved FIM methods overlook the instruction-following capability of LLMs, a crucial requirement in real-world code completion scenarios (Mozannar et al., 2024). In this work, we fill this gap by proposing IFIM, a training framework that preserves the infilling capability of FIM while significantly improving the model’s ability to follow instructions.
6.2. Instruction Tuning for Code LLMs
Instruction tuning, i.e., fine-tuning models on input-output pairs guided by natural language instructions, has proven to be a key technique for improving a model’s ability to follow user intent. This approach has been widely applied to general-purpose conversational LLMs such as DeepSeek-R1 (Guo et al., 2025) and Qwen 3 (Yang et al., 2025). Instruction tuning has also shown promise in enhancing code LLMs for generation tasks. Models like Magicoder (Wei et al., 2023) and WizardCoder (Luo et al., 2023), which are tuned on relatively small instruction datasets, demonstrate strong performance on code generation benchmarks. However, multiple studies (Ahmad et al., 2025; Gong et al., 2024; Pan et al., 2024) reveal that instruction-tuned models remain suboptimal for code completion tasks, when compared to FIM-pretrained base models. These findings highlight a gap in unifying instruction-following and infilling capabilities. Inspired by the effectiveness of instruction tuning in natural language domains, we argue that an effective instruction-tuning method tailored for code completion is needed.
7. Threats to Validity
7.1. Construct Validity
A potential threat lies in the quality of the instructions used for training and evaluation, as an LLM generated them and could suffer from hallucination. We mitigated this in several ways. For our large-scale training dataset, we followed established methodologies for synthetic data generation that have proven effective in similar studies. More importantly, for our evaluation benchmarks, where the quality of instructions is paramount for validating our claims, we implemented a rigorous manual review and revision process. This process, which resulted in an inter-rater agreement of 0.72, ensured that the instructions used to measure performance accurately reflected the desired code functionality, thus strengthening the validity of our conclusions.
7.2. Internal Validity
A threat to our study is that the observed performance gains might be attributable to the additional fine-tuning on more data, rather than the IFIM method itself. We explicitly mitigated this threat in our RQ3 experiment design by including a crucial control group: a 0% IFIM data ratio. In this configuration, the models were fine-tuned on the same code snippets from our dataset but without any of our synthesized instructions, keeping the data in the standard FIM format. This control group showed negligible performance gains over the original base models. Significant improvements in instruction-following were only achieved when the models were trained on data formatted with our IFIM method. This confirms that the performance boost comes directly from integrating instructions via the IFIM framework, not merely from exposure to more code, thereby ensuring the internal validity of our claims.
7.3. External Validity
Our experiments were conducted on a dataset predominantly composed of Python code (70.0%) and on models with up to 8 billion parameters. This could raise questions about generalizability to other programming languages and larger model architectures. We argue that focusing on Python, one of the most widely used languages, provides a strong and relevant test case for establishing the method’s feasibility. Furthermore, the IFIM method itself is language-agnostic. By demonstrating its success on two different state-of-the-art model families (Deepseek-Coder and Qwen2.5-Coder), we show that the approach is not tied to a specific architecture but offers a robust framework. This work serves as a crucial proof of concept, establishing a foundation that can be extended to other languages and larger models in future work.
8. Discussion
8.1. Alternative Data Source
As an initial work, our pipeline leverages LLM-generated synthetic instructions, a well-established approach that has demonstrated effectiveness in various studies (Wei et al., 2023; Luo et al., 2023; Huang et al., 2024). This method offers a controlled and scalable approach to generating high-quality training data. However, numerous alternative data sources remain underexplored, such as inline comments in source code, developer forum discussions, or real-world code completion logs. Such ”wild” data sources offer rich, diverse signals that could significantly enhance IFIM’s generalization capabilities. In the meantime, they are unstructured and noisy, requiring robust preprocessing and filtering techniques. Harnessing these resources effectively is a critical area for future research.
8.2. Seamless Instruction Handling in IDEs
In AI-assisted software development workflows, developers frequently encounter scenarios where they need to generate code that goes beyond simple autocompletion based on surrounding context. When developers require more sophisticated code generation that incorporates specific requirements, constraints, or implementation details, they face a fundamental choice in how to interact with AI assistance. In practice, developers have two primary approaches for obtaining instruction-aware code generation. The first approach involves describing the desired functionality within the current code context through comments or docstrings, relying on the code completion model to interpret these contextual cues. However, this method often falls short when dealing with complex requirements or when the base code language models lack sufficient instruction-following capabilities. The second approach involves switching to chat-based large language models to explicitly request help with code generation. While chat-based models typically demonstrate superior instruction comprehension and can engage in iterative refinement of requirements, the context-switching overhead disrupts the developer’s workflow, requiring them to leave their integrated development environment (IDE), manually transfer code context, and then integrate the generated code back into their project. IFIM provides an ideal solution where the LLM is specifically trained to interpret both implicit code context and explicit natural language instructions within a unified framework, eliminating the need for context switching while maintaining the instruction-following capabilities typically associated with chat-based models.
9. Conclusion and Future Work
In this paper, we addressed a critical limitation in modern code completion systems: the trade-off between a model’s ability to follow natural language instructions and its performance on fill-in-the-middle tasks. Existing instruction-tuning methods often degrade FIM capabilities, while standard FIM-trained models struggle to effectively use developer-provided instructions. To bridge this gap, we introduced Instruction-aware Fill-in-the-Middle (IFIM), a specialized tuning method that enhances FIM-trained models by extending their training objective to include an explicit instruction component, facilitated by a large-scale, model-generated dataset. Our evaluation on popular code LLMs demonstrates that IFIM significantly boosts instruction-following accuracy, improving the Pass@1 score from 84.6% to 93.6% on IHumanEval, without compromising the models’ original FIM performance. These results validate IFIM as a practical, backward-compatible, and effective solution for creating more helpful and responsive code completion systems.
For future work, we have identified several promising directions. First, we plan to expand our instruction dataset to encompass a wider variety of programming languages beyond the current Python-dominated collection to assess IFIM’s cross-lingual generalizability. Second, while our LLM-generated instructions proved effective, we intend to explore alternative and ”wild” data sources, such as inline comments, developer forum discussions, and real-world code completion logs, which could provide richer and more diverse training signals. This will also necessitate the development of robust filtering techniques to handle inherent noise. Finally, our experiments were constrained by available computational resources; a crucial next step will be to apply and evaluate the IFIM methodology on significantly larger, state-of-the-art code LLMs to understand its scaling properties.

### v2

@@
 Large Language Models (LLMs) have significantly advanced code completion, yet they often fail when the developer’s intent is underspecified in the code context.
-In practice, developers frequently add natural language instructions (e.g., comments) to clarify their intent.
-However, exisintg code LLMs applied for code completion systems merely undergoes a fill-in-the-middle (FIM) pre-training, which struggle to leverage this information effectively due to the lack of instruction-like training data.
+To address this, developers usually add natural language instructions (e.g., comments) into the code context to clarify their intent.
+However, existing code LLMs applied for code completion systems merely undergo a fill-in-the-middle (FIM) pre-training, which struggles to leverage this information effectively due to the lack of instruction-like training data.
 Existing instruction-tuning techniques, which improve instruction-following in general code generation, paradoxically degrade FIM performance, forcing a trade-off between instruction-following and infilling capabilities.
@@
 Moreover, this enhancement does not compromise the models’ original performance on FIM code completion tasks with no instructions provided.
-copyright: noneconference: 48th International Conference on Software Engineering; April 12–18, 2026; Rio De janeiro, Brazil
+copyright: none
 1. Introduction
-Modern Integrated Development Environments (IDEs) rely on code completion to accelerate development (Wang et al., 2023): given a partial program, a model predicts the missing sections and displays them to the programmer.
+Modern Integrated Development Environments rely on code completion to accelerate development (Wang et al., 2023): given a partial program, a model predicts the missing sections and displays them to the programmer.
 Large Language Models (LLMs) have pushed this feature from keyword-level suggestions to full-function generation, which adapts to different coding styles and languages with remarkable accuracy.
@@
 It thereby improves the model’s ability to generate contextually coherent and instruction-aligned infilling code.
-We evaluate IFIM on two code completion benchmarks IHumaneval and IRepoMasterEval, derived respectively from Humaneval-infilling (Fried et al., 2022) and RepoMasterEval (Wu et al., 2024) with code completion instructions added to their data samples.
+We evaluate IFIM on two code completion benchmarks IHumaneval and IRepoMasterEval, derived respectively from Humaneval-infilling (Fried et al., 2022) and RepoMasterEval (Wu et al., 2024) with human-annotated code completion instructions added to their data samples.
 To investigate the impact of IFIM training on widely used code completion LLMs, we experiment with two base code LLMs: Deepseek-Coder (DeepSeek-AI et al., 2024) and Qwen2.5-Coder (Hui et al., 2024).
-Our comprehensive experiments demonstrate that IFIM achieves remarkable improvements in instruction-following capability while maintaining the LLMs’ ability to complete the code without instructions.
+Our experiments demonstrate that IFIM achieves remarkable improvements in instruction-following capability while maintaining the LLMs’ ability to complete the code without instructions.
 Specifically, given the same instructions, our IFIM-trained variants can achieve better performance than the base model, with Pass@1 score averagely improved from 87.8% to 94.7% on IHumaneval and from 14.7% to 20.7% on IRepoMasterEval.
@@
 For example, on the IHumaneval benchmark, IFIM improves Deepseek-Coder from 68.6% to 78.2% and Qwen2.5-Coder from 76.0% to 76.3%.
-Moreover, we also conducted a series of ablation experiments to analyze the optimal position for inserting the instruction component, the ratio of IFIM processed data, and the advantages of using a standalone instruction component.
+Moreover, we also conducted a series of ablation experiments to analyze the optimal position for inserting the instruction component, the ratio of IFIM processed data, the advantages of using a standalone instruction component and the alignment between instructions and generated completions.
 These results position IFIM as a practical, backward-compatible upgrade path for existing LLM-based code completion systems.
@@
 Crucially, it achieves this while preserving or even improving performance when no instructions are given, validating IFIM as a practical and effective method for building more helpful code completion systems.
-Our source code and datasets are released at https://anonymous.4open.science/r/ifim-artifacts.
 2. Background and Motivation
@@
 2.1. Fill-in-the-Middle Code Completion
-The Fill-in-the-Middle (FIM) objective (Bavarian et al., 2022) was proposed to extend the conventional left-to-right, causal language-modeling paradigm with a crucial infilling capacity during pre-training.
+The FIM objective (Bavarian et al., 2022) was proposed to extend the conventional left-to-right, causal language-modeling paradigm with a crucial infilling capacity during pre-training.
 It grants the model the ability to generate a contiguous span of text that is missing from the middle of a document while conditioning on both the preceding (prefix) and the succeeding (suffix) context.
+Crucially, this new infilling capability does not compromise the model’s original performance in standard left-to-right generation.
 Code completion is the natural test-bed for this capability, as developers routinely need to insert new logic between already-written lines, such as the body of a newly declared function, missing import statements, or boilerplate loops.
@@
 However, the developer’s true, fine-grained intent while coding is a signal that is largely absent from the public code repositories used for pre-training.
-Much of that intent never reaches the committed source: it may surface as a fleeting search-engine query, a short Slack thread with a teammate, or even remain unspoken in the developer’s mind.
+Much of that intent never reaches the committed source code: it may surface as a fleeting search-engine query, a short Slack thread with a teammate, or even remain unspoken in the developer’s mind.
 The comments that are finally documented in the code are, therefore, an insufficient proxy for instructions.
@@
 This capability is cultivated through an instruction tuning stage, where models learn to follow human instructions and dialog formats, often through supervised learning on instruction-response pairs (Wei et al., 2021).
-However, the datasets used for this purpose, such as Magicoder (Wei et al., 2023) and Wizardcoder (Luo et al., 2023), primarily emphasize generating complete, assistant-style outputs rather than context-aware code insertions.
+However, the datasets used for this purpose (Wang et al., 2024; Wei et al., 2023; Luo et al., 2023) primarily emphasize generating complete, assistant-style outputs rather than context-aware code insertions.
 This focus comes at a cost: the training process largely overlooks the strict contextual and logical consistency required by FIM code completion, where new code must integrate seamlessly with its surrounding prefix and suffix.
@@
 An IFIM-trained LLM enables a developer to guide code completion by embedding instructions directly within their code.
+Notably, IFIM does not require developers typing instructions for every single completion they want.
+Instead, it serves as an effective solution to mitigate the unhelpful code completions caused by the under-specified developer intent.
 Since plain text would violate the programming language’s syntax, we propose writing instructions as specially marked inline comments.
@@
   • RQ3: What is the impact of varying the proportion of IFIM-processed data in the instruction tuning dataset?
-  • RQ4: Does representing instructions as a dedicated component outperform using instructions as inline comments?
+  • RQ4: Is representing instructions as a dedicated component effective?
+  • RQ5: To what extent does the code generated by IFIM-trained models align with completion instructions?
 4.1. Instruction Dataset
@@
 In our experiments, we apply IFIM training to the base versions since they are the LLMs used for code completion.
+For a comprehensive comparison, both the original base models and their instruction-tuned counterparts serve as baselines in our evaluation.
   • Deepseek-Coder (6.7B): Deepseek-Coder is a specialized code language model designed to excel in FIM tasks. It is commonly adopted as the baseline for various studies.
@@
 The final curated instruction benchmarks are respectively named as IHumanEval and IRepoMasterEval (IRME), which are used in the experiments.
+We also publicly release these two derived benchmarks to facilitate future research.
 4.4. Implementation Details
@@
 In this RQ, we systematically evaluate the effectiveness of IFIM, where we aim to assess two aspects: (1) whether IFIM can improve the instruction-following ability of the LLMs, and (2) whether IFIM can maintain the code completion ability when no instruction is given.
-Specifically, for the two base LLMs, Deepseek-coder and Qwen2.5-coder, we conduct IFIM training and obtain corresponding IFIM variants.
-The obtained models are evaluated with the two benchmarks, IHumanEval and IRME, under two settings: (1) with instruction and (2) without instruction.
-In the foIRMEr setting, we pair each completion request with its corresponding instruction, where the instructions for base models are written as a comment appended to the end of the prefix, while the instructions for IFIM variants are constructed following the IFIM mode.
-In the latter setting, we use pure prefixes and suffixes without any modification.
+For each base LLM (Deepseek-coder and Qwen2.5-coder), we compare its performance against two variants: (1) their official instruction-tuned version (instruct) and (2) our IFIM-trained version (ifim).
+The performance of these models are evaluated with the two benchmarks, IHumanEval and IRME, under two settings: (1) inputs with instructions and (2) inputs without instructions.
+In the “with instructions” setting, we pair each completion request with its corresponding instruction.
+For the base models, instructions are written as a comment appended to the end of the prefix.
+For the IFIM variants, instructions are constructed following the IFIM format.
+For the instruct variants, we embedded the instruction within a custom prompt designed for the fill-in-the-middle task.
+[^1] The prompt is available in our artifacts.
+In the ”without instructions” setting, inputs for the base and IFIM models are prefixes and suffixes without modification.
+For the instruct variants, however, we use the same prompt that asks the model to perform the fill-in-the-middle task, but without providing the instruction.
 The Pass@1 score is computed to assess the quality of the completion.
@@
 Setting Model Version Mode IHumanEval IRME
-w/ ins. Deepseek-Coder base PMS 84.6% 10.9%
-IFIM PIMS 93.6% 21.1%
-Qwen2.5-Coder base PSM 91.0% 18.4%
-IFIM PSIM 95.8% 20.3%
-w/o ins. Deepseek-Coder base PMS 68.6% 7.4%
-IFIM PIMS 78.2% 16.0%
-Qwen2.5-Coder base PSM 76.0% 10.2%
-IFIM PSIM 76.3% 13.3%
+Pass@1 Pass@1
+Inputs with instructions Deepseek-Coder base PMS 84.6% 10.90%
+instruct prompt 39.4% 2.10%
+ifim PIMS 93.6% 21.10%
+Qwen2.5-Coder base PSM 91.0% 18.40%
+instruct prompt 21.8% 3.60%
+ifim PSIM 95.8% 20.30%
+Inputs without instructions Deepseek-Coder base PMS 68.6% 7.40%
+instruct prompt 29.8% 0.60%
+ifim PIMS 78.2% 16.00%
+Qwen2.5-Coder base PSM 76.0% 10.20%
+instruct prompt 20.8% 1.90%
+ifim PSIM 76.3% 13.30%
 As reported in Table 1, the IFIM variants consistently outperform the base models when instructions are provided, across all the models and benchmarks.
@@
 Given the above results, we conclude that IFIM is an effective and feasible method for code LLMs to enhance their overall performance.
+Moreover, the performance of instruct models aligns with the results in previous studies (Ahmad et al., 2025; Gong et al., 2024; Pan et al., 2024), i.e., existing instruction tuning methods impair their capabilities of fill-in-the-middle code completion.
+This is evident in both settings.
+When instructions are provided via a prompt, the instruct models perform poorly, with scores dropping to as low as 39.4% (Deepseek-Coder) and 21.8% (Qwen2.5-Coder) on IHumanEval.
+The degradation is also severe in the infilling setting without instructions.
+For instance, Deepseek-Coder-instruct’s score drops from a base of 68.6% to 29.8% on IHumanEval and from 7.4% to a mere 0.6% on IRME.
 Answer to RQ1: Training with IFIM substantially improves instruction-following ability.
@@
 Notably, the I-before-M modes (i.e., placing the Instruction component immediately before the Middle component) consistently achieve the highest instruction-following performance compared with other IFIM modes.
-For Deepseek-Coder, the PIMS mode is the top perfoIRMEr on both IHumanEval (93.6%) and IRME (21.1%).
+For Deepseek-Coder, the PIMS mode is the top performer on both IHumanEval (93.6%) and IRME (21.1%).
 Similarly, for Qwen2.5-Coder, the PSIM mode achieves the best scores on both IHumanEval (95%) and IRME (20.3%).
@@
 5.4. RQ4: Impact of Standalone Instruction Component
-In this RQ, we seek to investigate the impact of our newly introduced Instruction component, compared with simply using instruction as inline comments.
+Compared to treating instructions as a separate, standalone component within the format, a simpler, more straightforward solution would be to train LLMs directly using instructions embedded as in-line comments.
+In this RQ, we seek to investigate the impact of training on this newly introduced Instruction component, compared with simply learning instruction through inline comments.
 Specifically, we derive two training datasets from our instruction dataset: one processed using the IFIM format and the other by appending the instruction as an inline comment to the prefix.
@@
 This confirms a hypothesis: LLMs can be confused when instructional text is mixed with code as comments, negatively impacting their primary code completion capabilities.
-The stark contrast between the significant gains from IFIM and the performance degradation from CFIM suggests that the structural separation of instructions from the code, as implemented in IFIM, is essential for the model to properly disambiguate, understand, and benefit from the provided guidance.
+The stark contrast between the significant gains from IFIM and the performance degradation from CFIM suggests that the structural separation of instructions from the code, as implemented in IFIM, is important for the model to properly disambiguate, understand, and benefit from the provided guidance.
 Answer to RQ4: The structural separation of instructions from prefix and suffix better facilitates LLMs’ understanding of natural language instructions during training.
+5.5. RQ5: Alignment between Instructions and Code Completions
+The IHumanEval and IRME benchmarks feature natural language instructions that are consistent with the surrounding code context.
+It could be questionable that the superior Pass@1 achieved by IFIM may come from the enhanced context understanding capability instead of the enhanced instruction-following capability.
+To this end, this RQ pushes the investigation further by understanding whether IFIM-trained models prioritize the natural language instruction over the implicit logic of the surrounding code.
+Specifically, we experiment to see when the intentions in the instruction and the code context are conflict, which side will the model follow.
+To create this conflict, we employed GPT-4 to systematically rewrite the instructions for each problem in the IHumanEval dataset.
+The rewriting process was guided by a specific prompt [^2] This prompt is available in our artifacts.
+that instructed GPT-4 to craft a new instruction that was deliberately different from the original one but still runnable in the code context.
+The written instructions are thus conflicting with the correct code expected by the code context.
+Some examples of these conflicting instructions are available in Figure 4.
+Using these newly created conflicting instructions for IHumanEval, we evaluate both the base and IFIM versions of DeepSeek-Coder.
+Our goal is to evaluate instruction alignment, i.e., whether the generated code align with the provided instruction.
+Therefore, in addition to Pass@1, we perform an human annotation to judge the alignment between the generated code and the completion instructions.
+Specifically, following a similar protocol to our benchmark creation, two authors independently classified each code completion based on whether the code implements the logic from the instruction.
+Notably, the annotators are required to only consider the alignment of the code’s logic with the instruction’s intent, irrespective of the code’s functional correctness.
+The evaluators achieved an inter-rater agreement of 0.69, and all disagreements were resolved through discussion to reach a final consensus.
+Table 5. Evaluating Instruction Alignment on Conflicting Instructions
+Model Version IHumanEval (Conflicting Instructions)
+Pass@1 Alignment
+Deepseek-Coder base 84.6% $\rightarrow$ 59.6% 54.4%
+ifim 93.6% $\rightarrow$ 56.7% 80.9%
+We report the pass@1 and the ratio of examples that are annotated as aligned in Table 5.
+The most direct measure of this is the alignment score.
+The IFIM-trained Deepseek-Coder achieves an alignment of 80.9%, indicating that in a vast majority of cases, it correctly generated code that implemented the logic of the new conflicting instructions.
+In contrast, the base model only achieved 54.4% alignment.
+This suggests that the base model was often confused by the conflict, frequently ignoring the commented instruction and instead generating code based on the original function signature and context.
+This conclusion is further reinforced by the Pass@1 scores.
+On the surface, both models show a significant drop in performance against the original test cases.
+This is an expected outcome, as correctly following a conflicting instruction should logically lead to a failure on the original tests However, the drop is more pronounced for the IFIM model (from 93.6% to 56.7%) than for the base model (from 84.6% to 59.6%).
+This steeper decline for the IFIM model is a direct consequence of its superior instruction-following ability.
+By successfully implementing the new instructions more often (as confirmed by the 80.9% alignment score), it naturally failed the now-irrelevant original tests at a higher rate.
+We further qualitatively demonstrate the instruction-following behaviors between base and IFIM models using five real examples, shown in Figure 4 In each case, the base model tends to ignore the commented instruction, defaulting to the logic suggested by the original function context.
+For instance, in Example 5 (HumanEval/63/L4), given the conflicting instruction that requires to check if n equals to 3, the base model still generates ‘if n == 2:‘, the code expected by the code context.
+Conversely, the IFIM model adheres to the new instruction, producing ‘if n == 2:‘.
+These examples provide qualitative evidence that IFIM training genuinely enhances the model’s ability to prioritize and follow explicit natural language instructions, even when they contradict the surrounding code context.
+Figure 4.
+A visual comparison of outputs produced by base and IFIM DeepSeek-Coder given conflicting instructions.
+The checkmark () indicates that the generated code aligns with the intent of the Conflicting Instruction, while the cross mark () indicates that it does not.
+Example 1: HumanEval/143/L3 Original Instruction: Check if the length of the word is 1.
+Conflicting Instruction: Check if the word starts with ’a’.
+Base model: if word[0] != ’a’: IFIM model: if word[0] == ’a’:
+Example 2: HumanEval/140/L12 Original Instruction: Append the current non-space character.
+Conflicting Instruction: Append an underscore character.
+Base model: new_text += text[i] IFIM model: new_text += "_"
+Example 3: HumanEval/3/L4 Original Instruction: Check if the balance is less than zero.
+Conflicting Instruction: Check if the balance is greater than zero.
+Base model: if balance < 0: IFIM model: if balance > 0:
+Example 4: HumanEval/13/L0 Original Instruction: Repeat the loop as long as b is not zero.
+Conflicting Instruction: Repeat the loop as long as a is not zero.
+Base model: while b != 0: IFIM model: while a != 0:
+Example 5: HumanEval/63/L4 Original Instruction: Check if n equals to 2.
+Conflicting Instruction: Check if n equals to 3.
+Base model: if n == 2: IFIM model: if n == 3:
+Answer to RQ5: IFIM-trained models demonstrate strong instruction alignment, prioritizing explicit instructions over the intention form code context.
+The IFIM variant followed conflicting instructions in 80.9% of cases, a significant increase from the base model’s 54.4%, confirming that the performance gains are attributable to enhanced instruction-following capability.
 6. Related Work
@@
 This will also necessitate the development of robust filtering techniques to handle inherent noise.
-Finally, our experiments were constrained by available computational resources; a crucial next step will be to apply and evaluate the IFIM methodology on significantly larger, state-of-the-art code LLMs to understand its scaling properties.
+Finally, our experiments were constrained by available computational resources; an important next step will be to apply and evaluate the IFIM methodology on significantly larger, state-of-the-art code LLMs to understand its scaling properties.
+Data-Availability Statement
+Our source code and datasets are released at https://anonymous.4open.science/r/ifim-artifacts.